OpenStudy (anonymous):

The perimeter of a triangle is 24x+35 . One side is 8x-9 and the other side is 12x+20 . Find the missing side.

jimthompson5910 (jim_thompson5910):

Let A, B and C be the three side lengths So A+B+C = perimeter, which means A+B+C = P

jimthompson5910 (jim_thompson5910):

The perimeter is 24x+35, so P = 24x+35 One side (say side A) is 8x-9 while the other (say side B) is 12x+20 So we go from A+B+C = P to (8x-9)+(12x+20)+C = 24x+35

jimthompson5910 (jim_thompson5910):

To find the third missing side, all you need to do is solve for C